OpenStudy (anonymous):

The members of an organization are 10 men and 11 women. What is the probability that a committee of three (selected at random) is all female?

OpenStudy (anonymous):

11/21

OpenStudy (anonymous):

But the committee is made up of three people out of the 21 people. So only three people are going to be picked.

OpenStudy (anonymous):

\[\frac{11}{21}\cdot\frac{10}{20}\cdot\frac{9}{19}\]

OpenStudy (anonymous):

Thank you SithsAndGiggles